# Graph Transformers: A Survey

### v1

@@ III-B2 Local Attention Mechanisms
Local attention mechanisms determine how each node computes its attention weights over a subset of nodes in the graph [85]. Local attention mechanisms can be broadly categorized into two types: message-passing attention mechanisms and spectral attention mechanisms.
Message-passing Attention Mechanisms. Message-passing attention mechanisms are built on the foundational framework of message-passing neural networks (MPNNs) [86], which iteratively aggregate messages from neighboring nodes to compute node representations. To enhance MPNNs, message-passing attention mechanisms employ self-attention to compute messages or aggregation weights [61, 87, 88, 39]. The computational complexity of this method is $O(E)$, where $E$ represents the number of edges in the graph.
One of the pioneering works in introducing message-passing attention mechanisms for graph transformers is GraphSAGE [89]. GraphSAGE proposes the use of mean pooling as an aggregation function and self-attention as a combination function. Additionally, GraphSAGE incorporates node sampling techniques to efficiently handle large-scale graphs. The function can be mathematically defined as:
@@
[…]
Spectral Attention Mechanisms. Spectral attention mechanisms are founded on the concept of transforming node features into a spectral domain, where the graph structure is encoded by eigenvalues and eigenvectors of the graph Laplacian matrix [91]. Spectral attention mechanisms employ self-attention to calculate spectral coefficients and spectral filters [92]. The computational complexity of this approach is $O(N)$.
Wang et al. [22] proposed graph isomorphism networks (GIN) as a spectral attention mechanism for graph transformers, employing an expressive aggregation function of sum pooling and self-attention. GIN also integrates a unique graph readout scheme via set functions to encapsulate global graph characteristics. The equations of GIN are as follows:
@@
Here, $\text{MLP}^{(l)}$ is a multi-layer perceptron, $\epsilon^{(l)}$ is a learnable parameter, READOUT indicates a set function that aggregates node output vectors into a graph output vector. $\mathbf{z}_{G}$ is the output vector of graph $G$. In addition, Nguyen et al. [93] introduced UGformer, a self-attention-based method for the spectral coefficient computation of each node using eigenvalues and eigenvectors of the Laplacian graph matrix. UGformer further integrates edge features into spectral computation via bilinear transformations.
Spectral attention mechanisms possess the ability to incorporate the structural information of a graph into the spectral domain, thereby offering potential benefits for certain tasks or domains. However, they are also accompanied by certain limitations, including high computational cost, memory consumption and sensitivity to graph size and density.
@@ IV Taxonomy of Graph Transformers
The past few years have witnessed a surge of interest in graph transformers. This section dives into four key categories dominating the current literature: shallow, deep, scalable, and pre-trained graph transformers. By analyzing representative models within each category, we aim to establish valuable guidelines for designing effective graph transformers.
@@ IV-A Shallow Graph Transformers
Shallow graph transformers represent a class of GNNs that leverage the power of self-attention to acquire node representations from data structured in graphs. Inspired by transformer models, which effectively capture long-range dependencies in sequential data through self-attention, shallow graph transformers extend this concept to graph data by computing self-attention weights based on both node features and graph topology [94]. The primary objective of shallow graph transformers is to achieve exceptional performance while minimizing computational complexity and memory usage.
Shallow graph transformers can be seen as a generalization of graph attention networks (GAT) [42]. GAT use a multi-head attention mechanism to calculate node embeddings. However, GAT has some limitations, such as the inability to model edge features and lack of diversity among attention heads [95]. Several GAT extensions have been proposed in the literature to address these issues. For example, GTN by Yun et al. [36] introduces edge-wise self-attention to incorporate edge information into node embeddings. Ahmad et al.[63] proposed the graph attention transformer encoder (GATE), which applies a masked self-attention mechanism to learn different attention patterns for different nodes. GATE also uses a position-wise feed-forward network and dropout to enhance model capacity and generalization. The summary of shallow graph transformer methods is given in Table II.
Shallow graph transformers are efficient and adaptable capable of handling various graph learning tasks and different types of graphs, but their lack of depth and recurrence may limit their ability to capture complex dependencies. Their performance can also be influenced by the choice of mask matrix and the number of attention heads, indicating a need for further research on their optimal design and regularization.
@@
Model Graph inductive bias Graph attention mechanisms Application Dataset Code availability
@@
PiT [96] No No No Yes (quadratic) No Image classification, object detection ImageNet, COCO, ImageNet-A Yes
WDPruning [97] No No No Yes (quadratic) No Image classification and image retrieval CIFAR-10 and ILSVRC-12 No
SVT [98] No No No Yes (quadratic) No Image classification Synthetic data and CIFAR-10 No
@@
Tree Transformer [38] Yes (local) No No Yes (quadratic) No Program translation and semantic parsing For2Lam, CoffeeScript-JavaScript, JOBS, GEO, ATIS Yes
@@ IV-B Deep Graph Transformers
Deep graph transformers consist of multiple self-attention layers stacked on top of each other, with optional skip connections, residual connections or dense connections between layers [102]. They are designed to achieve higher performance with increased model depth and complexity [103]. Deep graph transformers extend shallow graph transformers by applying self-attention layers to node features and graph topology hierarchically.
However, deep graph transformers also face several challenges that need to be addressed. One challenge is the difficulty of training deeper models, which can be mitigated by employing techniques, such as PairNorm introduced in DeeperGCN [104]. Another challenge is the over-smoothing problem, which can be addressed by using a gated residual connection and a generalized convolution operator as proposed in DeeperGCN. Additionally, the disappearance of global attention capacity and the lack of diversity among attention heads are challenges that can be tackled by approaches like DeepGraph [94]. DeepGraph incorporates substructure tokens and local attention to improve the focus and diversity of global attention.
Deep graph transformers, while complex, can achieve top-tier results on various graph learning tasks and adapt to different types of graphs and domains. However, their high computational cost, difficulty in optimization, and sensitivity to hyperparameters necessitate further research for optimal design and training. Table III summarizes the methods of deep graph transformers.
@@
HG-Transformer [102] No No No Yes (linear) No Multi-label text classification RCV1, RCV1-2K, AmazonCat-13K No
Unfolded Transformer [103] No No No Yes (quadratic) No Graph neural networks and NLP IMDB and SST2 Yes
@@
Tree Transformer [48] Yes (global and local) No No Yes (quadratic) Yes (message passing) Code representation learning Python150k and JavaScript150k Yes
MagLapNet [49] Yes (local and global) No No Yes (quadratic) Yes (spectral) Graph learning tasks such as correctness testing of sorting networks and source code understanding Synthetic graphs, OGB Code2 Yes
SAT [52] Yes (local and global) No No Yes (quadratic) Yes (message passing) Graph representation learning for graph and node property prediction ZINC, CLUSTER, PATTERN, OGBG-PPA, OGBG-CODE2 Yes
Edge Transformer [107] Yes Yes No Yes (quadratic) No Improve systematic generalization in natural language understanding and relational reasoning tasks CLUTRR, CFQ and COGS Yes
@@
KMCA [79] No No No Yes (quadratic) Yes (message passing) Time series analysis Synthetic, FordA, uWave, IMDB, ECG No
ABT-MPNN [87] No No Yes (pre-processing) Yes (quadratic and linear) No Molecular property prediction and biological activity prediction Johnson et al. [108], Tox21, ClinTox, ToxCast, HIV, QM8, ESOL, FreeSolv, Lipophilicity and QM8. Yes
UGformer [93] No No No Yes (quadratic) No Graph classification and text classification COLLAB, IMDB-B, IMDB-M, DD, PROTEINS, MUTAG, PTC, MR, R8, R52, Ohsumed Yes
@@ IV-C Scalable Graph Transformers
Scalable graph transformers are a category of graph transformers that tackle the challenges of scalability and efficiency when applying self-attention to large-scale graphs [114, 39, 53, 115]. These transformers are specifically designed to reduce computational cost and memory usage while maintaining or improving performance. To achieve this, various techniques are employed to reduce the complexity of self-attention, such as sparse attention, local attention, and low-rank approximation [12, 115]. Scalable graph transformers can be regarded as an enhancement of deep graph transformers addressing challenges, such as over-smoothing and limited capacity of global attention.
Several scalable graph transformer models have been proposed to enhance the scalability and efficiency of graph transformers. For instance, Rampášek et al. [39] introduced GPS, use low-rank matrix approximations to reduce computational complexity, and achieve state-of-the-art results on diverse benchmarks. GPS decouples local real-edge aggregation from a fully-connected transformer and incorporates different positional and structural encodings to capture graph topology. It also offers a modular framework that supports multiple encoding types and mechanisms for local and global attention. Cong et al. [116] developed DyFormer, a dynamic graph transformer that utilizes substructure tokens and local attention to enhance the focus and diversity of global attention. DyFormer employs a temporal union graph structure and a subgraph-based node sampling strategy for efficient and scalable training.
Scalable graph transformers are an innovative and efficient category of graph transformers that excel in handling large-scale graphs while minimizing computational cost and memory usage. However, scalable graph transformers face certain limitations, including the trade-off between scalability and expressiveness, the challenge of selecting optimal hyperparameters and encodings, and the absence of theoretical analysis regarding their convergence and stability. Consequently, further investigation is required to explore optimal designs and evaluations of scalable graph transformers for various applications. For a comprehensive overview of scalable graph transformer methods, please refer to Table IV.
@@
DIFFORMER [115] No No No Yes (linear) Yes (spectral) Node classification, image/text classification, spatial-temporal dynamics prediction Cora, Citeseer, Pubmed, ogbn-Proteins, Pokec, CIFAR, STL, 20News Yes
@@
Graph ViT/MLP-Mixer [117] Yes No No No Yes (message passing) Graph representation learning ZINC, MNIST, CIFAR10, MolTOX21, MolHIV, Peptide-func, Peptide-struct, CSL, EXP, SR25, LRGB, TreeNeighbourMatch Yes
@@
EXPHORMER [119] Yes (local and global) Yes (local and global) Yes (interleaving) Yes (linear) Yes (spectral) Graph learning and representation tasks CIFAR10, MalNet-Tiny, MNIST, CLUSTER, PATTERN, PascalVOC-SP, COCO-SP, PCQM-Contact, ogbn-arxiv, Yes
EGT [120] Yes (global) Yes (global) No Yes (quadratic) No Node and edge classification, graph classification and regression, transfer learning. PATTERN, CLUSTER, TSP, MNIST, CIFAR10, ZINC, PCQM4M, PCQM4Mv2, MolPCBA and MolHIV. Yes
@@
Higher-Order Transformers [123] No No No Yes (quadratic and linear) Yes (message passing) Graph and hypergraph modeling Synthetic chains, PCQM4M-LSC, Jets, Delaunay, Hyperedge prediction Yes
EGAT [73] No No Yes (interleaving) Yes (linear) No Graph representation learning and molecular property prediction AMLSim, Cora, Citeseer, PubMed No
L-CAT [90] No No No Yes (linear) No Node classification CSBM, Amazon, Facebook, GitHub, TwitchEN, Coauthor Physics, OGB (arxiv, products, mag, proteins) Yes
@@
Geo-ER [125] No No No Yes (linear) Yes (message passing) Geospatial Entity Resolution OSM, FSQ, Yelp Yes
GTA [126] No No No Yes (linear) Yes (message passing) Multivariate time series anomaly detection in IoT SWaT, WADI, SMAP, MSL Yes
@@
BSTG-Trans [132] Yes (local and global) Yes (local and global) No Yes (linear) Yes (message Passing) Long-term Pose Forecasting Human3.6M, HumanEva-I, Human360K Yes
@@
KG-Transformer [134] No No Yes (pre-processing) Yes (linear) No Text generation from knowledge graphs UMLS, PubMed, 2010 i2b2/VA, JNLPBA, BC5CDR, GAD, EU-ADR No
@@ IV-D Pre-trained Graph Transformers
Pre-trained graph transformers utilize large-scale unlabeled graphs to acquire transferable node embeddings [135]. These embeddings can be fine-tuned for downstream tasks with scarce labeled data that address the challenges of data scarcity and domain adaptation in graph learning tasks [136, 137]. These transformers are similar to pre-trained large language models (LLMs) and are trained on graph datasets using self-supervised learning objectives, such as masked node prediction [138], edge reconstruction [139], and graph contrastive learning [140]. These objectives aim to encapsulate the inherent properties of graph data independently of external labels or supervision [141]. The pre-trained model can be fine-tuned on a specific downstream task with a smaller or domain-specific graph dataset by incorporating a task-specific layer or loss function and optimizing it on labeled data [142]. This allows the pre-trained model to transfer the knowledge acquired from the large-scale graph dataset to the subsequent task, giving better performance compared to the training from scratch [142].
Pre-trained graph transformers face some challenges, such as the selection of appropriate pre-training tasks, domain knowledge incorporation, heterogeneous information integration, and pre-training quality evaluation [143]. To address these issues, KPGT [144] and KGTransformer [145] have been proposed. KPGT leverages additional domain knowledge for pre-training, while KGTransformer serves as a uniform Knowledge Representation and Fusion (KRF) module in diverse tasks. Despite their power and flexibility, pre-trained graph transformers encounter issues related to graph data heterogeneity and sparsity, domain adaptation, model generalization and performance interpretation. A summary of pre-trained graph transformer methods is provided in Table V.
@@
FormNetV2 [140] No No No Yes (linear) Yes (message passing) Form document information extraction FUNSD, CORD, SROIE No
@@
G-Adapter [148] No Yes (local) No Yes (quadratic) No Molecular graph tasks MolHIV, MolPCBA, FreeSolv, ESOL, BBBP, Estrogen-$\alpha$, Estrogen-$\beta$, MetStablow, MetStabhigh Yes
@@
Li et al. [149] Yes (local and global) No No Yes (linear) Yes (message passing) Movement synchrony estimation in autism therapy interventions PT13 and Human3.6MS No
Pellegrini et al. [150] No Yes No Yes (linear) No Disease prediction using graph convolutional networks TADPOLE and MIMIC-III Yes
Video Graph Transformer (VGT) [151] No No No Yes (linear) Yes (message passing) Video Question Answering (VideoQA) NExT-QA, TGIF-QA, MSRVTT-QA Yes
Tan et al. [152] No No No Yes (linear) No Few-shot node classification CoraFull, ogbn-arxiv, CiteSeer, Cora No
GPT-GNN [60] No Yes Yes No Yes Graph representation learning and graph mining OAG, Amazon, Reddit, OAG (citation) Yes
IV-E Design Guide for Effective Graph Transformers
Developing effective graph transformers requires meticulous attention to detail and careful consideration. This guide provides general principles and tips for designing graph transformers for various scenarios and tasks.
  • Choose the appropriate type of graph transformers based on the nature and complexity of your graph data and tasks. For simple and small graph data, a shallow graph transformer with a few layers may suffice. For complex and large graph data, a deep graph transformer with many layers can learn more expressive representations. For dynamic or streaming graph data, a scalable graph transformer is more efficient. Pre-trained graph transformers are more suitable for sparse or noisy graph data.
  • Design suitable structural and positional encodings for your graph data. These encodings capture the structure of graphs and are added to input node or edge features before feeding them to transformer layers. The choice of encodings depends on the characteristics of the graph data, such as directionality, weight, and homogeneity. The careful design of these encodings ensures their informativeness.
  • Optimize the self-attention mechanism for your graph data. Self-attention mechanisms compute attention scores among all pairs of nodes or edges in the graph, capturing long-range dependencies and interactions. However, it introduces challenges like computational complexity, memory consumption, overfitting, over-smoothing, and over-squashing. Techniques like sampling, sparsification, partitioning, hashing, masking, regularization, and normalization can be employed to address these challenges and improve the quality and efficiency of the self-attention mechanism.
  • Utilize pre-training techniques to enhance the performance of graph transformers. Pre-training techniques leverage pre-trained models or data from other domains or tasks transferring knowledge or parameters to a specific graph learning task. Methods like fine-tuning, distillation, and adaptation can be used to adapt pre-trained models or data. Utilizing pre-training techniques is particularly beneficial when a large amount of pre-training data or resources are available.

### v2

@@ III-B2 Local Attention Mechanisms
Local attention mechanisms determine how each node computes its attention weights over a subset of nodes in the graph [80]. These mechanisms can be broadly categorized into two types: message-passing attention mechanisms and spectral attention mechanisms.
Message-passing Attention. Message-passing attention mechanisms build on message-passing neural networks (MPNNs) [81], which iteratively aggregate messages from neighboring nodes to compute node representations. These mechanisms enhance MPNNs by using self-attention to compute messages or aggregation weights [56, 82, 83, 34]. The computational complexity is $O(E)$, where $E$ is the number of edges. GraphSAGE [84] introduced message-passing attention mechanisms for graph transformers, using mean pooling as an aggregation function and self-attention as a combination function, along with node sampling for large-scale graphs. The function is defined as:
@@
Here, $\mathbf{h}_{i}^{(l)}$ is the hidden vector of node $v_{i}$ at layer $l$, $\mathbf{W}^{(l)}$ is a linear transformation matrix, $\|$ is concatenation, and MEAN is mean pooling. $\alpha_{il}$ is the attention weight between node $v_{i}$ and layer $l$. Javaloy et al. [85] proposed L-CAT, using self-attention for both aggregation and combination, and incorporating edge features with bilinear transformations. Message-passing attention mechanisms effectively preserve and leverage local graph information but face scalability issues, including limited expressiveness and dependence on graph connectivity. Their ability to capture long-range dependencies and global graph information is also limited.
Spectral Attention. Spectral attention mechanisms transform node features into a spectral domain, encoding graph structure with eigenvalues and eigenvectors of the graph Laplacian matrix [86]. They use self-attention to calculate spectral coefficients and filters [87], with a computational complexity of $O(N)$. Wang et al. [22] proposed Graph Isomorphism Networks (GIN) as a spectral attention mechanism for graph transformers, utilizing sum pooling and self-attention. GIN integrates a graph readout scheme via set functions to encapsulate global graph characteristics, defined as:
@@
Here, $\text{MLP}^{(l)}$ is a multi-layer perceptron, $\epsilon^{(l)}$ is a learnable parameter, and READOUT aggregates node output vectors into a graph output vector $\mathbf{z}_{G}$. Nguyen et al. [88] introduced UGformer, which uses self-attention for spectral coefficient computation and incorporates edge features via bilinear transformations. Spectral attention mechanisms can incorporate graph structural information into the spectral domain, offering benefits for certain tasks. However, they face limitations such as high computational cost, memory consumption, and sensitivity to graph size and density.
@@ IV Taxonomy of Graph Transformers
Recent years have seen a surge of interest in graph transformers. This section examines four key categories in the literature—shallow, deep, scalable, and pre-trained graph transformers—organized to highlight distinctions between these categories. For example, shallow and deep graph transformers are compared based on model depth and ability to capture dependencies, while scalable and pre-trained transformers focus on handling large-scale graphs and leveraging unlabeled data, respectively. By expanding on specific examples of Graph Transformer models in each category, we aim to elucidate the architectural choices, strengths, and limitations that guide practitioners in selecting suitable approaches for their tasks. By analyzing representative models in each category, we aim to provide guidelines for designing effective graph transformers.
@@ IV-A Shallow Graph Transformers
Shallow graph transformers leverage self-attention to derive node representations from graph-structured data. Inspired by transformer models that capture long-range dependencies in sequential data, these transformers extend this concept to graph data by computing self-attention weights based on both node features and graph topology [89]. The primary goal of shallow graph transformers is to achieve high performance while minimizing computational complexity and memory usage. Unlike deep transformers, these models are optimized for efficiency and typically involve fewer layers, focusing on lightweight computation for smaller graphs or real-time applications.
These transformers can be viewed as a generalization of graph attention networks (GAT) [37], which use multi-head attention to compute node embeddings. However, GATs have limitations, such as the inability to model edge features and a lack of diversity among attention heads [90]. To address these issues, several extensions have been proposed. For instance, GTN by Yun et al. [31] introduces edge-wise self-attention to incorporate edge information into node embeddings. Similarly, Ahmad et al. [58] proposed the Graph Attention Transformer Encoder (GATE), which applies a masked self-attention mechanism to learn distinct attention patterns for different nodes. Such innovations demonstrate the adaptability of shallow graph transformers, but their limited depth constrains their ability to capture complex graph dependencies, which are better addressed by deeper models. In practice, shallow models often excel in lower-resource or time-sensitive scenarios due to their reduced computational overhead and simpler optimization. However, their smaller capacity may limit performance on highly complex tasks, making them most effective when graph structures are relatively simple or when inference speed is a priority. The summary of shallow graph transformer methods is given in Table II.
@@
Tree Transformer [33] Yes (local) No No Yes (linear) No Program translation and semantic parsing For2Lam, CoffeeScript-JavaScript, JOBS, GEO, ATIS Yes
@@ IV-B Deep Graph Transformers
Deep graph transformers consist of multiple stacked self-attention layers with optional skip, residual, or dense connections between layers [94]. They aim for higher performance through increased model depth and complexity [95, 96], applying self-attention layers hierarchically to node features and graph topology. Unlike shallow models, these transformers can capture complex, multi-level dependencies in large and heterogeneous graphs, making them suitable for tasks requiring deep feature extraction.
Training deeper models is challenging, often requiring techniques like PairNorm from DeeperGCN to mitigate difficulties [97]. Over-smoothing can be addressed using gated residual connections and generalized convolution operators, also proposed in DeeperGCN. Other challenges, such as the loss of global attention capacity and lack of diversity among attention heads, can be tackled by approaches like DeepGraph [89], which uses substructure tokens and local attention to enhance focus and diversity. Although deep graph transformers achieve state-of-the-art performance on many graph tasks, their optimization remains sensitive to hyperparameter tuning, computational costs, and architectural choices, which require further study. Deeper architectures generally exhibit stronger representational power, enabling them to model complex graph relationships and long-range dependencies. Nonetheless, this comes at a higher computational cost, greater risk of vanishing gradients, and potential overfitting. Successful deep transformer designs often incorporate specialized normalization and regularization layers to preserve training stability and manage computational overhead. Table III summarizes the methods of deep graph transformers.
@@
MagLapNet [44] Yes (local and global) No No Yes (quadratic) Yes (spectral) Source code understanding Synthetic graphs, OGB Code2 Yes
SAT [47] Yes (local and global) No No Yes (quadratic) Yes (message passing) Node property prediction ZINC, CLUSTER, PATTERN, OGBG-PPA, OGBG-CODE2 Yes
Edge Transformer [100] Yes Yes No Yes (quadratic) No Generalization in natural language understanding and relational reasoning tasks CLUTRR, CFQ and COGS Yes
@@
ABT-MPNN [82] No No Yes (pre-processing) Yes (quadratic and linear) No Molecular property prediction Johnson et al. [101], Tox21, ClinTox, ToxCast, HIV, QM8, ESOL, FreeSolv, Lipophilicity and QM8. Yes
UGformer [88] No No No Yes (quadratic) No Graph classification COLLAB, IMDB-B, IMDB-M, DD, PROTEINS, MUTAG, PTC, MR, R8, R52, Ohsumed Yes
@@ IV-C Scalable Graph Transformers
Scalable graph transformers address the challenges of scalability and efficiency in applying self-attention to large-scale graphs [107, 34, 48, 108]. These models reduce computational cost and memory usage while maintaining or improving performance by employing techniques such as sparse attention, local attention, and low-rank approximation [12, 108]. Compared to deep transformers, scalable models are optimized for handling large datasets and real-world applications with limited resources.
Several models have been proposed to enhance scalability and efficiency. Rampášek et al. [34] introduced GPS, which uses low-rank matrix approximations to reduce computational complexity and achieve state-of-the-art results on diverse benchmarks. GPS decouples local real-edge aggregation from a fully-connected transformer and incorporates various positional and structural encoding to capture graph topology, offering a modular framework for local and global attention mechanisms. Cong et al. [109] developed DyFormer, a dynamic graph transformer that utilizes substructure tokens and local attention to enhance global attention diversity. While scalable graph transformers address key issues of memory and computation, they often involve trade-offs between scalability and expressiveness, requiring further research on optimal encoding and task-specific adaptations. Scalable models typically strike a balance between performance and resource demands, employing approximations or partitioning methods to handle large datasets without exhaustive computation. However, aggressive approximation can lead to reduced expressiveness, and domain-specific strategies may be needed to recover performance in certain tasks or graph structures. For a comprehensive overview, see Table IV.
@@
DIFFORMER [108] No No No Yes (linear) Yes (spectral) Node classification Cora, Citeseer, Pubmed, ogbn-Proteins, Pokec, CIFAR, STL, 20News Yes
@@
EXPHORMER [111] Yes (local and global) Yes (local and global) Yes (interleaving) Yes (linear) Yes (spectral) Graph classification CIFAR10, MalNet-Tiny, MNIST, CLUSTER, PATTERN, PascalVOC-SP, COCO-SP, PCQM-Contact, ogbn-arxiv, Yes
EGT [112] Yes (global) Yes (global) No Yes (quadratic) No Node and edge classification PATTERN, CLUSTER, TSP, MNIST, CIFAR10, ZINC, PCQM4M, PCQM4Mv2, MolPCBA and MolHIV. Yes
@@
EGAT [68] No No Yes (interleaving) Yes (linear) No Molecular property prediction AMLSim, Cora, Citeseer, PubMed No
@@
BSTG-Trans [121] Yes (local and global) Yes (local and global) No Yes (linear) Yes (message Passing) Pose forecasting Human3.6M, HumanEva-I, Human360K Yes
@@
KG-Transformer [123] No No Yes (pre-processing) Yes (linear) No Text generation UMLS, PubMed, 2010 i2b2/VA, JNLPBA, BC5CDR, GAD, EU-ADR No
@@ IV-D Pre-trained Graph Transformers
Pre-trained graph transformers leverage large-scale unlabeled graphs to acquire transferable node embeddings, which can be fine-tuned for downstream tasks with limited labeled data [124, 125, 126]. These models are analogous to pre-trained language models, enabling knowledge transfer and reducing data dependence by using general-purpose embeddings. Similar to pre-trained large language models (LLMs), these transformers are trained using self-supervised learning objectives like masked node prediction [127], edge reconstruction [128], and graph contrastive learning [129], capturing the inherent properties of graph data without external labels [130]. The pre-trained model can then be fine-tuned on specific tasks using a smaller or domain-specific graph dataset, incorporating task-specific layers or loss functions, leading to better performance compared to training from scratch.
Challenges for pre-trained graph transformers include selecting appropriate pre-training tasks, incorporating domain knowledge, integrating heterogeneous information, and evaluating pre-training quality [131]. Models like KPGT [132], which uses additional domain knowledge, and KGTransformer [133], which acts as a uniform Knowledge Representation and Fusion (KRF) module, have been proposed to address these issues. While pre-trained models are versatile, their application is constrained by challenges related to domain-specific adaptations, evaluation consistency, and interpretability of embeddings. Further work is needed to standardize benchmarks and improve transfer learning capabilities. Compared to shallow or scalable transformers, pre-trained models often excel at adapting to new tasks with minimal labeled data, leveraging rich representations gleaned from large-scale training. However, this advantage can be limited if the pre-training domain diverges significantly from the target domain or if specialized graph structures require model re-parameterization. For a comprehensive overview, see Table V.
@@
G-Adapter [136] No Yes (local) No Yes (quadratic) No Molecular graph tasks MolHIV, MolPCBA, FreeSolv, ESOL, BBBP, $\alpha, Estrogen$-$\beta$, MetStablow, MetStabhigh Yes
@@
Li et al. [137] Yes (local and global) No No Yes (linear) Yes (message passing) Movement synchrony estimation PT13 and Human3.6MS No
Pellegrini et al. [138] No Yes No Yes (linear) No Disease prediction TADPOLE and MIMIC-III Yes
Video Graph Transformer (VGT) [139] No No No Yes (linear) Yes (message passing) Video Question Answering NExT-QA, TGIF-QA, MSRVTT-QA Yes
Tan et al. [140] No No No Yes (linear) No node classification CoraFull, ogbn-arxiv, CiteSeer, Cora No
GPT-GNN [55] No Yes Yes No Yes Graph representation learning OAG, Amazon, Reddit, OAG (citation) Yes